OpenStudy (anonymous):

find the area of the region. between the two curves. and the two equations are y = 7x^2 y^2 = (1/7)x

OpenStudy (anonymous):

Set them equal to find the boundaries a and b. Y1 = Y2 Set up an integral:\[\int\limits_{a}^{b} Y1+Y2\]

OpenStudy (anonymous):

ops, top equation - bottom equation in the integral

OpenStudy (anonymous):

\[\int\limits_{a}^{b}Ytop - Ybottom \]

OpenStudy (anonymous):

7x^2 = sqrt of (1/7)x

OpenStudy (anonymous):

should i square both sides to get rid of the radical?

OpenStudy (anonymous):

sure yu can do that